A bag contains 2 white marbles and 7 purple marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. What is the probability of selecting a purple marble and then a white marble? What is the probability of selecting two white marbles? Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row? Show your work.
The probability of selecting a purple marble and then a white marble is 0.19.
The probability of selecting two white marbles is 0.03.
The probability of picking two purple marbles is greater.
Probability is used to determine the likelihood that an event would occur.
The probability of selecting a purple marble and then a white marble = (number of purple marbles / total number of marbles) x (number of white marbles / total number of marbles- 1)
7/9 x 2/8 = 0.19
The probability of selecting two white marbles = (number of white marbles / total number of marbles) X (number of white marbles / total number of marbles- 1)
2/9 X 1/8 = 0.03
The probability of selecting two purple marbles = 7/9 x 6/8 = 0.58
To learn more about probability, please check: https://brainly.com/question/23722574
#SPJ5